OpenStudy (anonymous):

1.00 L of a gas at standard temperature and pressure is compressed to 473 mL. What is the new pressure of the gas? Please show steps

OpenStudy (anonymous):

as the question does not say anything about temp. change, we can safely assume that temp. is constant during the decrease in volume. \[P1*V1=P2*V2\] BOYLES LAW P1=1 ATM P2=? V1=I LIT V2=0.473LIT plug the values, P2 is unknown hence P2=(P1*V1)/V2 Ans is 2.11 atm